Use the expansion of (1+x)^n; to expand: 1/(1+7x)^4, in ascending powers of x, as far as x^3 term.
Answer by venugopalramana(3286)

Use the expansion of (1+x)^n; to expand: 1/(1+7x)^4, in ascending powers of x, as far as x^3 term.
1/(1+7X)^4=(1+7X)^(-4)
=1+[(-4)/1!](7X)+[(-4)(-5)/2!](7X)^2+[(-4)(-5)(-6)/3!](7X)^3+....
=1-28X+490X^2-6860X^3+...